OpenStudy (anonymous):

Solve by complete the square: y=-3x^2-9x-0.75

OpenStudy (anonymous):

step one is to divide by \(-3\)

OpenStudy (anonymous):

by "solve' do you mean "solve for \(x\) if \(-3x^2-9x-0.75=0\) "?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then divide by \(-3\) and start with \[x^2+3x+.25=0\]

OpenStudy (anonymous):

you need the leading coefficient two be 1 to complete the square

OpenStudy (anonymous):

then subtract to get \[x^2+3x=-\frac{1}{4}\] and go right to \[\left(x+\frac{3}{2}\right)^2=-\frac{1}{4}+\frac{9}{4}\] and work from there